OpenStudy (anonymous):

HELP ME PLEASE .57 + 5/8= the answer is supposed to be 1.195. but i'm not sure about the steps on how to get the answer?

OpenStudy (tygrr321):

this is pretty easy, now tell me how u got 1.195

OpenStudy (tygrr321):

tag me when u stop being afk @reticens

OpenStudy (anonymous):

\[\frac{57}{100}+\frac{5}{8}=\frac{239}{200}=1.195 \]